9. Hansen Delivery just paid an annual dividend of $2.50 per share. The company has been reducing the dividends by 4 percent annually and expects to continue doing so. What is the current market value for this company’s stock if your required rate of return is 13 percent?

$14.12 Please Show All Work.
Answer #1
Formula for Dividend growth model
Po = D1/(re-g)
where,
Po = Current market value
D1 =Dividend for next year = $2.5*0.96 = $2.4
G=growth rate = -4% or -0.04
re= required rate of return = 13% or 0.13
Po = 2.4/(0.13-(-0.04))
Po = 2.4/(0.17)
Po = Current market value = $14.12
Answer #3

Using Dividend Discount Model

Current Market Price = [Dividend Paid * (1 + Growth Rate)] / [Required Rate of Return - Growth Rate]

Current Market Price = [2.50 * (1 - 0.04)] / [13% - (-4%)]

Current Market Price = 2.40 / 17%

Current Market Price = $14.12
Answer #2

ANSWER :


Given :


D0 = 2.50 ($)

r = 13 % = 0.13

g = - 4 % = - 0.04


So,

D1 = D( 1 - 0.04) = 2.50 * 0.96 = 2.40 ($)


Now,


P= D1 / (r - g)  = 2.40 / (0.13 - (- 0.04)) = 14.12 ($)


So, Price of $14.12 can be paid for this stock (ANSWER)


answered by: Tulsiram Garg